OpenStudy (anonymous):

Determine the critical #s, the intervals on which the function is increasing or decreasing, all relative maximum points and relative minimum points, and the intervals of concavity. f(x)= (x^2+4) / (x-8)^2

OpenStudy (anonymous):

For which function ????

OpenStudy (anonymous):

oh sorry! f(x)= (x^2+4)/(x-8)^2

OpenStudy (anonymous):

Any work on this so far ?

OpenStudy (anonymous):

uh yea i did the derivative and got something like 2(x-x^2+4) but idk it ocudl be wrong

OpenStudy (mathmale):

Let's use Equation Editor to express your function.\[f(x)= (x^2+4)/(x-8)^2=\frac{ x^2+4 }{ (x-8)^2 }\] Don't have to do this, but seeing the function in this form makes it so much clearer, for me, anyway. This is a quotient function To find its derivative, apply the quotient rule. Are you familiar with that?

OpenStudy (anonymous):

Here to find derivative we need to use quotient rule of differentiation Are you familiar with it ???

OpenStudy (mathmale):

(Great minds think alike! :) )

OpenStudy (anonymous):

uh i think

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

what i did was the deonominator and the derivative of the numerator and then vice versa

OpenStudy (mathmale):

Suggest you make up a list of common derivative formulas and refer to it often. The Quotient Rule is \[\frac{ d }{ dx }\frac{ f }{ g }=\frac{ g*f ' - f*g }{ g^2 }\]

OpenStudy (anonymous):

yea thats wat i did

OpenStudy (anonymous):

the answer i got was 2x-2(x^2-4)/(x-8) and then i triedd simplifying it more

OpenStudy (mathmale):

then, replacing the functions f and g and their derivatives in this formula by their values in the given problem, \[\frac{ d }{ dx }\frac{ f }{ g }=\frac{ (x-8)^2*2x - (x^2+4)*2(x-8)^1 }{ ((x-8)^2)^2 }\]

OpenStudy (mathmale):

Please simplify this as much as you can.

OpenStudy (mathmale):

Just to be clear: f(x)=x^2+4 f' (x) = 2x g(x)=(x-8)^2 g'(x)=2(x-8)

OpenStudy (mathmale):

You need to find the "critical values" and then the corresponding "critical points." To do this, set the numerator of the derivative = to 0 and solve for x. Were I doing this, I would factor out (x-8) from the numerator first.

OpenStudy (anonymous):

yes thats wat i did

OpenStudy (anonymous):

i think. my answer was 2x-2(x^2-4)/(x-8)

OpenStudy (mathmale):

Please explain. Was this expression your derivative, the numerator of the derivative, the denominator, or what? It's really hard to communicate these expressions through ordinary typing; if you were to use Equation Editor or Draw, your intent would be clearer.

OpenStudy (anonymous):



OpenStudy (anonymous):

is that better?

OpenStudy (mathmale):